OpenStudy (anonymous):

a regular polygon has 29 sides what is the size of each interior angle

OpenStudy (campbell_st):

there are several ways to do this... 1. find the size of 1 exterior angle = 360/29 then use the fact that the interior angle is supplementary 2. find the angle sum of the polygon using the formula \[Sum = (n -2) \times 180\] where n = the number of sides then divide the angle sum by n to find 1 angle. Hope it helps

OpenStudy (anonymous):

thank you i think i understand now